FEL Upgrade Project Weekly Brief
October 2-6, 2000


We continued preparations for the Oct. 23rd project review and for the start of operations for the Fall User run next week.

WBS 3 (Beam Physics)

We continued to make progress on a cost-effective modification of the baseline lattice that can accommodate the addition of a chicane for the
UV wiggler/optical cavity.  The concept will be discussed at the weekly FEL Upgrade meeting on Monday (Oct. 9th).

WBS 4 (Injector)

We started preliminary design work on the option of putting a cathode shield and cesium channels into the ball cathode.  We are running
further tests on the N2 implanted electrode with a 6mm gap at 20MV/m, a long duration test is currently in progress.

At this week’s AVS Symposium in Boston we learned about additional data that CERN has gathered in-situ getter coatings that we are planning to
use in the upgraded gun.  CERN has quantified the surface characteristics and microstructure of the alloy (Ti-Zr-Vn) that activates at the lowest temperatures (200-200C).  Kimball Physics showed us a concept for a small Cs ion source that we may also consider for cathode cesiation.

The film thickness gage that will be used in the cesiator test stand has been released to us so we will modify it to remove the leak.  In the
interim, the cesium has been loaded into its tube and is ready for pump down.


We completed the technical discussions with the vendor chosen for the waveguide couplers.  The final contract can now be written.   Also this week we received the proposal from KEK (described in last week’s weekly brief) concerning KEK’s interest in electropolish treatments of the FEL module SRF

WBS 6 (RF)

All 8 klystrons were installed in zone 4.  Shorts were placed on 4 of the 8 waveguides.  The debug and test process is to start next week on
zone 4.  Installation efforts continued in zone 3.

Glassman High Voltage has agreed to provide us a slave unit to work with one of our existing drive units.  We will send them the multiplier and
measuring stacks along with the drive so they make any necessary changes.  They will integrate and test the system at their plant.  We
will make any final modifications after it is installed again here.  The requisition and sole source justification will be written next week.

WBS 8  (I & C)

The week was spent preparing for the hot checkout next week and the user run following. Cables for the Optical Beam Position Monitors (OBPM) have
been installed. The 32 channel VME board design/layout is complete and is in production.  The printed circuit boards will arrive Monday, these
should be stuffed and checked out by mid-week.  The commercial 64 channel A/D boards arrived this week, the software is ready to go -
there were only minor differences from what was already in use.  The design for the LiTaO3 detector preamp and board design is complete and
was sent out today.  The "beta" for the system should work down to 100 microseconds, more effort will make the optical diagnostics fully
operational down to 10 microseconds.  Again, thanks to the excellent software support the FEL has received an alternative positioning system
has been brought on line - this is based on the Spiricon Pyrocam.  By connecting the output to the MaxVideo VME video processor one can track
X & Y position.  This is operational in EPICS.

Other improvements to the user lab include picomotor controls for 32 additional steering motors, there will be cables and mux'd outputs for
up to 8 motors per lab. The user LAN has been wired for 4 ports per user lab. This will allow remote control of users data acquisition PCs from
the control room or any other lab and allow web access and email without an IP address.

The components for the vacuum controls and ion pump power supplies have been arriving, these are being stored in the FEL building prior to
shipping out to a local fab shop for assembly. The 3" BPM detector documentation is being sent out for bids.

WBS 9 (Transport)

o Change of Plans for prototyping and drawing packages:
1. The dipoles in the corners of the IR machine where the beam for the UV beam exits and re-enters have a new requirement imposed by the UV
beam optics.  They need to run at half field, making beam exit and enter at 21 degrees to the return leg.  The pole width needs to get wider at
that end of the magnet to maintain the 1 part per 10,000 field in the region of the angled beam.

2. The magnetic model has given us confidence that we can add shims to the edge of the poles and combined with the recommended pole width,
achieve the magnetic tolerance in this "H" style magnet.

3. The optical chicane and Pi Bends are planned to be window frame dipoles where we only achieved our tolerance in the IR Demo by the use
of a Purcell Gap and µ metal pole tip.  Since models of the Purcell gap never were able to benchmarked this field uniformity gain.  These are
the magnets that need prototyping.

o We will make all four GX magnets like the two for the UV beam to save time in contracting, fabrication, inventory and measurement.
o We will not prototype these GX magnets.
o We will draw up the GW magnet in house for prototyping with various higher field replacement metals for the µ metal.
o We will set aside the now obsolete GX drawing package as a template from which an outside designer could make the GQ and new GX drawing
o Poles:  Results from the CMM inspection of the First Article quadrant were reviewed with Dave McCay and the vendor.  The data showed a
systematic offset in the part.  The offset was traced to a poor choice of technique in how the part was fixtured on the CNC.  Several other
errors were reviewed with the vendor.  A partial piece will be machined out of aluminum and CMM'd to see if the vendor is capable of meeting the
tolerances.  Throughout all this the J-Lab Survey and Alignment Group has been most accommodating in performing several CMM inspections on
short notice and with quick turn around in the data.

o   Budgetary Quote: No word from Everson.  Greg Naumovich from Everson said last Friday that the engineer doing the quote has been out ill.

o Best and finals have been requested of the vendors for the engineering services contract.

WBS 10 (Wiggler)

Bids are due on the optical klystron dispersion section Monday.

WBS 11 (Optics)

The optics team has updated the Project Requirements Document for revisions to the optical cavity section which reflect our new data on
low loss optics.


This week we continued the cross-training of ten CEBAF operators on FEL operations.  We re-cesiated the photogun and high voltage conditioned
the gun for operation at 320 kV In preparation for next week’s start of the Fall User Run.  We will have “hot check-out” of the FEL on Oct. 9-11
and start user operations sometime on the 11th depending on our success with start-up.

Privacy and Security Notice